# On the Design of Expressive and Trainable Pulse-based Quantum Machine Learning Models

### v1

@@ 4 Numerical Simulations
where $\Theta=\{\theta_{1}(t),\cdots,\theta_{m+p}(t)\}$. The scaling factor $\theta_{0}$ is introduced to adapt the function represented by the model to the range of the dataset. Without loss of generality, the domain of x is always set as $\mathcal{X}=[-1,1]^{m}$ [22].
In the numerical simulations, the control pulses are chosen in a piecewise-constant form. Each control pulse is divided into $K$ sub-pulses, analogous to the number of layers in gate-based QNN models, with the sub-pulse amplitudes used to parameterize the model. In all simulations, the sampling period is set to $\Delta t=T/K=0.1$, and the model is trained by Adam optimizer [50] with a learning rate of $0.05$. The simulations are conducted using MATLAB on a computer equipped with two 10-core Xeon CPUs and 208 GB RAM.
4.1 Expressivity of pulse-based models with dynamical symmetry
For the convenience of model description, we adopt the following notation
$\sigma_{\alpha}^{(k)}=\mathbb{I}_{2}\otimes\cdots\otimes\sigma_{\alpha}\otimes\cdots\otimes\mathbb{I}_{2},\leavevmode\nobreak\ \leavevmode\nobreak\ \alpha=\mathrm{x},\mathrm{y},\mathrm{z},$
@@ 4.1 Expressivity of pulse-based models with dynamical symmetry
The system is not controllable as it possesses a six-dimensional dynamical symmetry algebra:
@@
We choose the observable $M=\sigma_{\rm z}^{(1)}\sigma_{\rm z}^{(2)}$ and it can be calculated that
@@
Figure 2: The approximation of the polynomial function (15) using the two-qubit pulse-based model (13) under $\mathfrak{so}(4)$ dynamic symmetry. The inset figure displays the trained control pulses.
To validate the above prediction, we apply the model (13) to approximate the following randomly selected target function:
@@
from which $200$ points are evenly sampled across $\mathcal{X}=[-1,1]$ to train the model with pulse duration $T=20$ (corresponding to $K=200$). The simulation results presented in Fig. 2 indicate that the trained quantum machine learning (QML) model accurately fit the target function when $|\psi_{0}\rangle=\left(\frac{2}{\sqrt{5}}|0\rangle+\frac{1}{\sqrt{5}}|1\rangle\right)\otimes|0\rangle$. However, when the initial state is $|\psi_{0}\rangle=|0\rangle\otimes|0\rangle$, the model can only approximate the even part of $f_{0}(x)$ due to the violation of the necessary condition in Theorem 1.
We also test the fitting of other randomly selected functions, which can all be approximated by the model, provided that the pulse duration is sufficiently long. The required pulse durations tend to be longer when the target function exhibits greater curvature or includes many higher-order power terms.
We further examine the approximation capability of pulse-based models for more complex bivariate functions using the following Hamiltonian:
@@
which is similar to (13) with an additional Hamiltonian encoding $\sigma_{\rm y}^{(1)}\sigma_{\rm y}^{(2)}$, and the observable is chosen as $M=\sigma_{\rm z}^{(1)}+\sigma_{\rm z}^{(2)}$. This model also exhibits $\mathfrak{so}(4)$ dynamic symmetry, and it can be computed that
@@
for all cases where $k_{1}+k_{2}$ is odd. Accordingly, the initial state is designated as $|\psi_{0}\rangle=\frac{1}{\sqrt{2}}|0\rangle\otimes\left(|0\rangle+|1\rangle\right)$, which can be verified to satisfy the necessary condition in Theorem 1.
@@
For comparison, we train two pulse-based models. In the first case, we set the control pulses $\theta_{1}(t)=\theta_{2}(t)\equiv 1$, which considerably limits the tunability of the coefficients $C_{k_{1}k_{2}}(\Theta)$. As illustrated in Fig. 3(a), the fitting progressively aligns with the surface of the target bivariate function as $T$ increases; however, the precision remains limited no matter how $T$ is prolonged. In the second case, we allow $\theta_{1}(t)$ and $\theta_{2}(t)$ to vary to enhance the tunability of coefficients $C_{k_{1}k_{2}}(\Theta)$, thereby releasing more expressivity compared to the first case. The fitting results in Fig. 3(b) validate this expectation, as it nearly perfectly matches the target surface when $T=4$, with a fitting error of $L(\Theta)\leq 10^{-3}$.
Figure 3: The approximation results of bivariate function by two-qubit models with (a) $\theta_{1}(t)=\theta_{2}(t)=1$ and (b) $\theta_{1}(t)$ and $\theta_{2}(t)$ being freely tunable.
4.2 The balance between expressivity and trainability
The simulations above demonstrate that pulse-based models can be expressive when uncontrollable due to dynamic symmetry. This establishes a framework for the design of large-scale pulse-based QML models that are both expressive and trainable. Subsequently, we select four classes of multi-qubit pulse-based models that exhibit various dynamic symmetries.
Model 1. The first model class possesses ${\mathfrak{su}}(2)$ dynamic symmetry with the following model Hamiltonian:
@@ 4.2 The balance between expressivity and trainability
is the $2^{n}$-dimensional unitary irreducible representation of ${\mathfrak{su}}(2)$ (see the detailed description in appendix). In this case, the observable is chosen as $M=\frac{2}{2^{n}-1}J_{\rm z}$ and the initial state is chosen as $|\psi_{0}\rangle=|0\rangle^{\otimes n}$.
Model 2. The second model class possesses ${\mathfrak{su}}(2)^{\oplus n}$ dynamic symmetry with the following non-interacting $n$-qubit system:
@@
with $2n$ control pulses. The observable is selected as $M=\frac{1}{n}\sum_{k=1}^{n}\sigma_{\rm z}^{(k)}$, and the initial state $|\psi_{0}\rangle=|0\rangle^{\otimes n}$.
Model 3. The third model class is selected from Ref. [51] that possesses ${\mathfrak{so}}(n)$ dynamic symmetry. The Hamiltonian of the model is as follows:
@@
Model 4. The fourth model class is a fully controllable model utilizing the following circularly coupled $n$-qubit system in [22]:
@@
$=$ $\leavevmode\nobreak\ x\sum_{k=1}^{n}\sigma_{\rm z}^{(k)}+\sum_{k=1}^{n}\left[\theta_{x}^{(k)}(t)\sigma_{\rm x}^{(k)}+\theta_{y}^{(k)}(t)\sigma_{\rm y}^{(k)}\right]$
@@
with $3n$ control pulses. The observable is selected as $M=\sigma_{\rm z}^{(1)}$ and the initial state as $|\psi_{0}\rangle=|0\rangle^{\otimes n}$.
@@
The initial states and measurement observables selected in the four models all satisfy the necessary condition in Theorem 1 for expressivity. In addition, the observables are selected from the associated Lie algebra $\mathfrak{g}$ for the sake of trainability [41]. The theoretical variance of the measurement output (5) is presented in Tab. 1 (see the appendix for detailed calculations), which indicates that the variance is inversely proportional to the dimension of the dynamic Lie algebra $\mathfrak{g}$.
@@
During the numerical simulations, these models were trained to fit the univariate function (15). Due to the limitation of computational resource, we use up to six qubits. As can be seen in the inset figure of Fig. 4(a), each model can approximate the target function with arbitrarily small error when the pulse duration $T$ is sufficiently extended. This indicates that the model expressivity grows with $T$, and thus provides a measure to quantify the degree of expressivity. Here, we adopt the minimal pulse duration required for the training loss to fall below $10^{-3}$ to evaluate the expressivity (see the inset for illustration).
@@
Based on this measure, the simulation results in Fig. 4(a) compare the variation of expressivity with respect to system size across the four model classes. It is observed that, except the special case of $\mathfrak{su}(2)$ whose dimensionality is constant, the expressivity can be enhanced by increasing the number of qubits. Furthermore, when the number of qubits is sufficiently large, expressive power increases in the order of $\mathfrak{su}(2)$, $\mathfrak{su}(2)^{\otimes n}$, $\mathfrak{so}(n)$, and $\mathfrak{su}(2^{n})$, implying that expressivity can be improved by enlarging the dynamical symmetry Lie algebra.
@@
Regarding the trainability of the models, we estimate the variance of the loss function $L(\Theta)$ in the numerical simulation. Note that the variance of $L(\Theta)$ is different from ${\rm Var}[f(\mathbf{x},\Theta)]$ listed in Table. 1 as $L(\Theta)$ is a quadratic function of $f(\mathbf{x},\Theta)]$, and an analytic formula is unavailable. Therefore, we randomly pick 1000 sets of control pulses to numerically evaluate $L(\Theta)$ and its variance. Since the variance is also dependent on the pulse duration $T$, we make $T$ sufficiently long so that the estimated variance gradually converges to a stationary value, corresponding to the condition under which the model approaches a $2$-design.
@@
Figure 4(b) illustrates the relation between the stationary variance and the number of qubits. Consistent with the theoretical analysis in Tab. 1, the variance related to $\mathfrak{su}(2^{n})$ exhibits an exponential decline as the number of qubits $n$ increases, indicating the presence of a barren plateau. In contrast, the variance associated with the $\mathfrak{su}(2)$ symmetry remains constant owing to its constant dimensionality, while it decreases at a polynomial rate for the $\mathfrak{su}(2)^{\otimes n}$ and $\mathfrak{so}(n)$ symmetries.
@@
These observations suggest that barren plateaus can be effectively avoided by restricting the dynamical evolution to submanifolds within the quantum state Hilbert space through engineered dynamic symmetry, thereby achieving a good balance between expressivity and trainability. It is noted that the overall expressivity and trainability are highly dependent on the selection of the Hamilton structure, the initial state and the observable. For instance, the expressivity of the larger $\mathfrak{so}(n)$ models is inferior to that of the smaller $\mathfrak{su}(2)^{\otimes n}$ models when the number of qubits is small, as the latter has more tunable control pulses. Nonetheless, their rates of decrease align with predictions, indicating that $\mathfrak{so}(n)$ models will become more expressive than $\mathfrak{su}(2)^{\otimes n}$ models when $n$ is sufficiently large.
@@
Table 1: The symmetry Lie algebra, dimension and the variance of the inference output (5) for each model.
 Model  $\mathfrak{g}$  ${\rm dim}(\mathfrak{g})\leavevmode\nobreak\$ ${\rm Var}\left[f(\textbf{x},\Theta)\right]$
@@ 4 ${\mathfrak{su}}(2^{n})$ $4^{n}-1$ $\frac{1}{2^{n}+1}$
Figure 4: The comparison of expressivity and trainability between model classes under $\mathfrak{su}(2)$, $\mathfrak{su}(2)^{\otimes n}$, $\mathfrak{so}(n)$ and $\mathfrak{su}(2^{n})$ dynamic symmetries: (a) The expressivity (evaluated by the minimal pulse duration when the training loss reaches $10^{-3}$) versus qubit number for each example . (b) The sample variance of loss function versus qubit number for each example. The inset shows how the minimal pulse duration is obtained. (20).
5 CONCLUSION AND OUTLOOK
In conclusion, we established a comprehensive framework for designing practical pulse-based QML models that are both expressive and trainable. The design incorporates Dyson polynomial-series expansion alongside existing Lie algebraic theories related to the barren plateau phenomenon in quantum systems exhibiting dynamic symmetry. Theoretical analysis and numerical simulations demonstrate the potential to leverage dynamic symmetry to construct expressive and trainable pulse-based models suitable for hardware-efficient deployment on NISQ devices.
It remains an open question whether the proposed necessary condition is sufficient for the expressivity of QML models with dynamic symmetry. Simulation results indicate that complete expressivity is not always attainable for models that satisfy this condition, particularly when control pulses are constrained. We conjecture that sufficiency is achieved when all control pulses are unrestricted and the pulse duration is sufficiently long, but this requires verification in the future.
We also note that, even if the pulse-based models are expressive and trainable, the loss landscape may contain undesirable local traps induced by dynamic symmetry [52], which can attract the training process to suboptimal models. Although this issue is rarely encountered in our simulations, the possibility remains, and thus the training algorithm needs to be carefully designed.
Besides the trade-off between model expressivity and trainability, an important issue for future investigation is their trade-off between model generalizability, which pertains to the model’s performance on unseen data. The overall balance among these three attributes holds significant practical importance and will be the focus of subsequent studies.
Appendix A Validation of expressivity and variance calculation for the four examples

### v2

@@ IV NUMERICAL SIMULATIONS
where $\Theta=\{\theta_{1}(t),\cdots,\theta_{m+p}(t)\}$. The scaling factor $\theta_{0}$ adapts the function represented by the model to the range of the dataset. Without loss of generality, the domain of x is set as $\mathcal{X}=[-1,1]^{m}$ [22].
In the numerical simulations, the control pulses are chosen in piecewise-constant form. Each control pulse is divided into $K$ sub-pulses, analogous to the number of layers in gate-based QNN models, with the sub-pulse amplitudes used to parameterize the model. The sampling period is set to $\Delta t=T/K=0.1$, and the model is trained using the Adam optimizer [50] with a learning rate of $0.05$. The simulations are conducted using MATLAB with the QuTiP package [qutip] for quantum dynamics simulation on a computer equipped with two 10-core Xeon CPUs and 208 GB RAM.
IV.1 Expressivity of pulse-based models with dynamical symmetry
For the convenience of model description, we adopt the notation
$\sigma_{\alpha}^{(k)}=\mathbb{I}_{2}\otimes\cdots\otimes\sigma_{\alpha}\otimes\cdots\otimes\mathbb{I}_{2},\penalty 10000\ \penalty 10000\ \alpha=\mathrm{x},\mathrm{y},\mathrm{z},$
@@ IV.1 Expressivity of pulse-based models with dynamical symmetry
The system is not controllable because it possesses a six-dimensional dynamical symmetry algebra:
@@
We choose the observable $M=\sigma_{\rm z}^{(1)}\sigma_{\rm z}^{(2)}$ and calculate that
@@
Figure 2: Approximation of the polynomial function (15) using the two-qubit pulse-based model (13) under $\mathfrak{so}(4)$ dynamic symmetry. The inset displays the trained control pulses.
To validate this prediction, we apply model (13) to approximate the randomly selected target function:
@@
from which $200$ points are evenly sampled across $\mathcal{X}=[-1,1]$ to train the model with pulse duration $T=20$ (corresponding to $K=200$). The simulation results presented in Fig. 2 indicate that the trained QML model accurately fits the target function when $|\psi_{0}\rangle=\left(\frac{2}{\sqrt{5}}|0\rangle+\frac{1}{\sqrt{5}}|1\rangle\right)\otimes|0\rangle$. However, when the initial state is $|\psi_{0}\rangle=|0\rangle\otimes|0\rangle$, the model can only approximate the even part of $f_{0}(x)$ due to violation of the necessary condition in Theorem 1.
We also tested the fitting of other randomly selected functions, which can all be approximated by the model, provided that the pulse duration is sufficiently long. The required pulse durations tend to be longer when the target function exhibits greater curvature or includes many higher-order power terms.
We further examine the approximation capability of pulse-based models for more complex bivariate functions using the Hamiltonian:
@@
which is similar to (13) with an additional Hamiltonian encoding $\sigma_{\rm y}^{(1)}\sigma_{\rm y}^{(2)}$, and the observable is chosen as $M=\sigma_{\rm z}^{(1)}+\sigma_{\rm z}^{(2)}$. This model also exhibits $\mathfrak{so}(4)$ dynamic symmetry, and we compute that
@@
for all cases where $k_{1}+k_{2}$ is odd. Accordingly, the initial state is designated as $|\psi_{0}\rangle=\frac{1}{\sqrt{2}}|0\rangle\otimes\left(|0\rangle+|1\rangle\right)$, which satisfies the necessary condition in Theorem 1.
@@
For comparison, we train two pulse-based models. In the first case, we set the control pulses $\theta_{1}(t)=\theta_{2}(t)\equiv 1$, which considerably limits the tunability of the coefficients $C_{k_{1}k_{2}}(\Theta)$. As illustrated in Fig. 3(a), the fitting progressively aligns with the surface of the target bivariate function as $T$ increases; however, the precision remains limited regardless of how long $T$ is extended. In the second case, we allow $\theta_{1}(t)$ and $\theta_{2}(t)$ to vary to enhance the tunability of coefficients $C_{k_{1}k_{2}}(\Theta)$, thereby increasing expressivity compared to the first case. The fitting results in Fig. 3(b) confirm this expectation, as they nearly perfectly match the target surface when $T=4$, with a fitting error of $L(\Theta)\leq 10^{-3}$.
Figure 3: Approximation results of bivariate function by two-qubit models with (a) $\theta_{1}(t)=\theta_{2}(t)=1$ and (b) $\theta_{1}(t)$ and $\theta_{2}(t)$ freely tunable.
IV.2 Balance between expressivity and trainability
These simulations demonstrate that pulse-based models can be expressive when uncontrollable due to dynamic symmetry. This establishes a framework for designing large-scale pulse-based QML models that are both expressive and trainable. We select four classes of multi-qubit pulse-based models exhibiting various dynamic symmetries.
Model 1. The first model class exhibits ${\mathfrak{su}}(2)$ dynamic symmetry with Hamiltonian:
@@ IV.2 Balance between expressivity and trainability
is the $2^{n}$-dimensional unitary irreducible representation of ${\mathfrak{su}}(2)$ (see Appendix for detailed description). The observable is chosen as $M=\frac{2}{2^{n}-1}J_{\rm z}$ and the initial state as $|\psi_{0}\rangle=|0\rangle^{\otimes n}$.
Model 2. The second model class exhibits ${\mathfrak{su}}(2)^{\oplus n}$ dynamic symmetry with a non-interacting $n$-qubit Hamiltonian:
@@
with $2n$ control pulses. The observable is $M=\frac{1}{n}\sum_{k=1}^{n}\sigma_{\rm z}^{(k)}$, and the initial state $|\psi_{0}\rangle=|0\rangle^{\otimes n}$.
Model 3. The third model class, selected from Ref. [51], exhibits ${\mathfrak{so}}(n)$ dynamic symmetry with Hamiltonian:
@@
Model 4. The fourth model class is fully controllable, utilizing the circularly coupled $n$-qubit system from [22]:
@@
$=$ $\penalty 10000\ x\sum_{k=1}^{n}\sigma_{\rm z}^{(k)}+\sum_{k=1}^{n}\left[\theta_{x}^{(k)}(t)\sigma_{\rm x}^{(k)}+\theta_{y}^{(k)}(t)\sigma_{\rm y}^{(k)}\right]$
@@
with $3n$ control pulses. The observable is $M=\sigma_{\rm z}^{(1)}$ and the initial state $|\psi_{0}\rangle=|0\rangle^{\otimes n}$.
@@
The initial states and measurement observables in the four models all satisfy the necessary condition in Theorem 1 for expressivity. Additionally, the observables are selected from the associated Lie algebra $\mathfrak{g}$ to ensure trainability [41]. The theoretical variance of the measurement output (5) is presented in Table 1 (see Appendix for detailed calculations), indicating that the variance is inversely proportional to the dimension of the dynamic Lie algebra $\mathfrak{g}$.
@@
We trained these models to fit the univariate function (15). Due to computational resource limitations, we used up to six qubits. As shown in the inset of Fig. 4(a), each model approximates the target function with arbitrarily small error when the pulse duration $T$ is sufficiently extended. This indicates that model expressivity grows with $T$, providing a measure to quantify expressivity. We adopt the minimal pulse duration required for the training loss to fall below $10^{-3}$ as a measure of expressivity (see inset for illustration).
@@
Using this measure, Fig. 4(a) compares how expressivity varies with system size across the four model classes. We observe that, except for $\mathfrak{su}(2)$ whose dimensionality is constant, expressivity increases with the number of qubits. Moreover, for sufficiently large numbers of qubits, expressive power increases in the order $\mathfrak{su}(2)<\mathfrak{su}(2)^{\otimes n}<\mathfrak{so}(n)<\mathfrak{su}(2^{n})$, indicating that expressivity improves with larger dynamical symmetry Lie algebras.
@@
Regarding model trainability, we estimate the variance of the loss function $L(\Theta)$ numerically. Note that the variance of $L(\Theta)$ differs from ${\rm Var}[f(\mathbf{x},\Theta)]$ in Table 1 since $L(\Theta)$ is a quadratic function of $f(\mathbf{x},\Theta)]$, and no analytic formula is available. Therefore, we randomly select 1000 sets of control pulses to evaluate $L(\Theta)$ and its variance numerically. Since the variance also depends on the pulse duration $T$, we choose $T$ sufficiently long for the estimated variance to converge to a stationary value, corresponding to the model approaching a $2$-design.
@@
Fig. 4(b) shows the relationship between the stationary variance and the number of qubits. Consistent with the theoretical analysis in Table 1, the variance for $\mathfrak{su}(2^{n})$ decreases exponentially with increasing qubit number $n$, indicating a barren plateau. In contrast, the variance for $\mathfrak{su}(2)$ remains constant due to its constant dimensionality, while decreasing polynomially for $\mathfrak{su}(2)^{\otimes n}$ and $\mathfrak{so}(n)$.
@@
These results demonstrate that barren plateaus can be avoided by restricting dynamical evolution to submanifolds of the quantum state Hilbert space through engineered dynamic symmetry, achieving a favorable balance between expressivity and trainability. We note that overall expressivity and trainability depend strongly on the Hamiltonian structure, initial state, and observable. For example, $\mathfrak{so}(n)$ models with larger $n$ are less expressive than smaller $\mathfrak{su}(2)^{\otimes n}$ models for small qubit numbers, since the latter have more tunable control pulses. However, their scaling behavior matches theoretical predictions, indicating that $\mathfrak{so}(n)$ models become more expressive than $\mathfrak{su}(2)^{\otimes n}$ models for sufficiently large $n$.
@@
Table 1: Symmetry Lie algebra, dimension, and variance of the inference output (5) for each model.
Model $\mathfrak{g}$ ${\rm dim}(\mathfrak{g})$ ${\rm Var}\left[f(\textbf{x},\Theta)\right]$
@@ 4 ${\mathfrak{su}}(2^{n})$ $4^{n}-1$ $\frac{1}{2^{n}+1}$
Figure 4: Comparison of expressivity and trainability between model classes under $\mathfrak{su}(2)$, $\mathfrak{su}(2)^{\otimes n}$, $\mathfrak{so}(n)$ and $\mathfrak{su}(2^{n})$ dynamic symmetries: (a) Expressivity (evaluated by minimal pulse duration when training loss reaches $10^{-3}$) versus qubit number. (b) Sample variance of loss function versus qubit number. The inset shows how the minimal pulse duration is obtained.
V CONCLUSION AND OUTLOOK
In conclusion, we have established a comprehensive framework for designing practical pulse-based QML models that are both expressive and trainable. Our main contributions include: (i) a necessary condition for expressivity in terms of the initial state, observable, and dynamical Lie algebra; (ii) numerical validation of the condition and its impact on expressivity; and (iii) a demonstration of the balance between expressivity and trainability in various models with dynamic symmetry. Our approach combines Dyson polynomial-series expansion with existing Lie algebraic theories for barren plateaus in quantum systems with dynamic symmetry. Theoretical analysis and numerical simulations demonstrate that dynamic symmetry enables construction of expressive and trainable pulse-based models suitable for hardware-efficient implementation on NISQ devices.
Whether the proposed necessary condition is sufficient for expressivity of QML models with dynamic symmetry remains an open question. Our simulations indicate that complete expressivity is not always achieved in models satisfying this condition, particularly with constrained control pulses. We conjecture that sufficiency holds when all control pulses are tunable and the pulse duration is sufficiently long, but this requires future verification.
We also note that even if pulse-based models are expressive and trainable, the loss landscape may contain undesirable local minima due to dynamic symmetry, which can trap the training process in suboptimal solutions. Although we rarely encountered this issue in our simulations, this possibility necessitates careful design of training algorithms.
Beyond the trade-off between model expressivity and trainability, an important direction for future work is the trade-off with model generalizability, which concerns model performance on unseen data. Balancing these three attributes has significant practical importance and will be the focus of future work.
Appendix A VALIDATION OF EXPRESSIVITY AND VARIANCE CALCULATIONS FOR THE FOUR MODELS